(iv) The taxi charges in a city consist of a fixed charge together with the charge for the distance covered. For a distance of 10 km, the charge paid is Rs 105 and for a journey of 15 km, the charge paid is Rs 155. What are the fixed charges and the charge per km? How much does a person have to pay for travelling a distance of 25 km?

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Solution: Let the fixed charge be Rs x and per km charge be Rs y. According to the question, x + 10y = 105 …………….. (1) x + 15y = 155 …………….. (2) From (1), we get x = 105 – 10y ………………. (3) Substituting the value of x in (2), we get 105 – 10y + 15y = 155 5y = 50 y = 10 …………….. (4) Putting the value of y in (3), we get x = 105 – 10 × 10 = 5